# A topological counting rule for shells

### v1

@@ Abstract
The results presented here establish a macroscopic counting rule for macroscopic isometric deformations of simply connected shells. By combining the static–geometric analogy with the Hill–Mandel lemma, the spaces of macroscopic isometries and of admissible macroscopic stresses were proven to be complementary and isomorphic. Their dimensionality then yields precisely three macroscopic isometric modes, just as it yields precisely three admissible macroscopic stresses. The same arguments reveal an exact algebraic relationship between the effective elasticity moduli providing a microstructure-independent constraint on the homogenized constitutive law. Note that the fact that the shell’s surface was taken to be a smooth graph is a matter of convenience; The same techniques can be adapted to handle jumps in the tangent plane (e.g., straight and curved creases) as well as overhangs; see, e.g., the classical tools used in [28] as well as the measure theory formalism used in [44, 45]. By contrast, simple connectedness is an essential, both sufficient and necessary, ingredient of the theory. Indeed, topology is known to severely impact structural stiffness, as it does for instance in Saint-Venant’s problem of torsion for simply or multiply connected sections; the present context is no exception.

### v2

@@ Conclusion
The results presented here establish a counting rule for effective isometric deformations in simply connected shells. By combining the static–geometric analogy with the Hill–Mandel lemma, the spaces of effective isometries and of admissible effective stresses were proven to be isomorphic, orthogonal and Lagrangian (relative to a symplectic form). Their dimensionality then yields precisely three effective isometric modes, just as it yields precisely three admissible effective stresses. The same arguments reveal an exact algebraic relationship between the effective elasticity moduli providing a microstructure-independent constraint on the homogenized constitutive law. Note that the fact that the shell’s surface was taken to be a smooth graph is a matter of convenience. By contrast, simple connectedness is an essential, both sufficient and necessary, ingredient of the theory. Indeed, topology is known to severely impact structural stiffness, as it does for instance in Saint-Venant’s problem of torsion for simply or multiply connected sections; the present context is no exception.
@@ Acknowledgements.
Supplemental Material: Proof of the Hill-Mandel lemma (14)
Consider a field of membrane stresses $\bm{\sigma}$ that are statically admissible in the sense of (1). Let $\bm{\varepsilon}$ be the field of membrane strains of an ansatz $(\mathbf{u},w)$ satisfying (4), i.e.,
$\begin{bmatrix}u_{\alpha}\\ w\end{bmatrix}=\begin{bmatrix}E_{\alpha\beta}x_{\beta}+z\chi_{\alpha\beta}x_{\beta}\\ -\frac{1}{2}\chi_{\mu\nu}x_{\mu}x_{\nu}\end{bmatrix}+\begin{bmatrix}\tilde{u}_{\alpha}\\ \tilde{w}\end{bmatrix}$ (23)
where the periodic correction has been denoted with tilde. Thus, the membrane strain is
$\bm{\varepsilon}=\mathbf{E}+z\bm{\chi}+\bm{\nabla}^{s}\tilde{\mathbf{u}}+\bm{\nabla}\tilde{w}\otimes^{s}\bm{\nabla}z.$ (24)
Now expand
$\left\langle{\bm{\sigma}:\bm{\varepsilon}}\right\rangle-\left\langle{\bm{\sigma}}\right\rangle:\mathbf{E}-\left\langle{z\bm{\sigma}}\right\rangle:\bm{\chi}$ (25)
$=$ $\left\langle{\sigma_{\alpha\beta}(\tilde{u}_{\alpha,\beta}+\tilde{w}_{\beta}z_{\alpha})}\right\rangle,$ (by symmetry of $\bm{\sigma}$)
$=$ $\left\langle{(\sigma_{\alpha\beta}(\tilde{u}_{\alpha}+\tilde{w}z_{\alpha}))_{\beta}}\right\rangle-\left\langle{\sigma_{\alpha\beta,\beta}(\tilde{u}_{\alpha}+\tilde{w}z_{\alpha})+\sigma_{\alpha\beta}z_{\alpha\beta}\tilde{w}}\right\rangle$ (by the product rule)
$=$  $\,0,$ (by periodicity & static admissibility)
to get the desired lemma. It is crucial to note that the lemma holds for virtual fields, i.e., regardless of whether $\bm{\sigma}=\mathbf{c}:\bm{\varepsilon}$ or not.
Supplemental Material: Proof of equation (10)
Begin by proving the polarization identity
$\begin{bmatrix}\mathbf{E}_{1}\\ \bm{\chi}_{1}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}_{2}\\ \bm{\chi}_{2}\end{bmatrix}=\left\langle{\bm{\varepsilon}_{1}:\mathbf{c}:\bm{\varepsilon}_{2}}\right\rangle,$ (26)
for any $(\mathbf{E}_{i},\bm{\chi}_{i})$ where $\bm{\varepsilon}_{i}$ minimizes the corresponding average strain energy density. Indeed, for any such strains, $\bm{\varepsilon}_{3}\equiv\bm{\varepsilon}_{1}+\bm{\varepsilon}_{2}$ minimizes the strain energy density for prescribed $(\mathbf{E}_{3}\equiv\mathbf{E}_{1}+\mathbf{E}_{2},\bm{\chi}_{3}\equiv\bm{\chi}_{1}+\bm{\chi}_{2})$ by linearity. Thus,
$\begin{bmatrix}\mathbf{E}_{i}\\ \bm{\chi}_{i}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}_{i}\\ \bm{\chi}_{i}\end{bmatrix}=\left\langle{\bm{\varepsilon}_{i}:\mathbf{c}:\bm{\varepsilon}_{i}}\right\rangle,\quad(i=1,2,3).$ (27)
Then, expand
$\left\langle{\bm{\varepsilon}_{3}:\mathbf{c}:\bm{\varepsilon}_{3}}\right\rangle-\left\langle{\bm{\varepsilon}_{1}:\mathbf{c}:\bm{\varepsilon}_{1}}\right\rangle-\left\langle{\bm{\varepsilon}_{2}:\mathbf{c}:\bm{\varepsilon}_{2}}\right\rangle$ (28)
and simplify thanks to the symmetry of $\mathbf{c}$ and $\mathbf{A}$ to get the polarization identity, as desired.
Now to prove (10), proceed by proving a double inclusion for each equality, starting with strains.
  1. 1.
    Let $(\mathbf{E}_{o},\bm{\chi}_{o})\in\text{Ker}\mathbf{A}$ and let $\bm{\varepsilon}_{o}$ be the corresponding membrane strain. Then, by definition $\left\langle{\bm{\varepsilon}_{o}:\mathbf{c}:\bm{\varepsilon}_{o}}\right\rangle=0$ and $\bm{\varepsilon}_{o}=\mathbf{0}$ by the positive definiteness of the membrane stiffness $\mathbf{c}$. Hence, $(\mathbf{E}_{o},\bm{\chi}_{o})$ is the effective strain of an admissible isometric deflection.
  2. 2.
    Let $(\mathbf{E}_{o},\bm{\chi}_{o})$ be the effective strain of an admissible isometric deflection, i.e., with $\bm{\varepsilon}_{o}=\mathbf{0}$. Then, by the polarization identity
    $\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}_{o}\\ \bm{\chi}_{o}\end{bmatrix}=\left\langle{\bm{\varepsilon}:\mathbf{c}:\bm{\varepsilon}_{o}}\right\rangle=0,$ (29)
    for all $(\mathbf{E},\bm{\chi})$ with corresponding $\bm{\varepsilon}$. Thus,
    $\mathbf{A}\begin{bmatrix}\mathbf{E}_{o}\\ \bm{\chi}_{o}\end{bmatrix}=0,\quad\text{and}\quad(\mathbf{E}_{o},\bm{\chi}_{o})\in\text{Ker}\mathbf{A}.$ (30)
  3. 3.
    Let $(\bm{\Sigma},\mathbf{M})\in\text{Im}\mathbf{A}$. Then there exists $(\mathbf{E},\bm{\chi})$ such that
    $\begin{bmatrix}\bm{\Sigma}\\ \mathbf{M}\end{bmatrix}=\mathbf{A}\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}$. (31)
[…]
    $\begin{bmatrix}\hat{\mathbf{E}}\\ \hat{\bm{\chi}}\end{bmatrix}^{T}\begin{bmatrix}\bm{\Sigma}\\ \mathbf{M}\end{bmatrix}=\left\langle{\hat{\bm{\varepsilon}}:\bm{\sigma}}\right\rangle=\hat{\mathbf{E}}:\left\langle{\bm{\sigma}}\right\rangle+\hat{\bm{\chi}}:\left\langle{z\bm{\sigma}}\right\rangle$ (32)
[…]
    $\bm{\Sigma}=\left\langle{\bm{\sigma}}\right\rangle,\quad\mathbf{M}=\left\langle{z\bm{\sigma}}\right\rangle.$ (33)
    Hence, any $(\bm{\Sigma},\mathbf{M})$ in the image of $\mathbf{A}$ is an admissible effective state of stress.
  4. 4.
    Let $(\bm{\Sigma},\mathbf{M})$ be an admissible effective stress obtained by averaging an admissible periodic membrane stress $\bm{\sigma}$. Then, by the Hill-Mandel lemma,
    $\bm{\Sigma}:\mathbf{E}_{o}+\mathbf{M}:\bm{\chi}_{o}=0,$ (34)
    for all $(\mathbf{E}_{o},\bm{\chi}_{o})\in\text{Ker}\mathbf{A}$. Then, $(\bm{\Sigma},\mathbf{M})\in(\text{Ker}\mathbf{A})^{\perp}=\text{Im}\mathbf{A}$, by symmetry of $\mathbf{A}$.
This concludes the proof that
$\begin{split}\text{Ker}\mathbf{A}&=\{(\left\langle{z\mathbf{H}_{w}}\right\rangle,-\left\langle{\mathbf{H}_{w}}\right\rangle)\mid w\text{ is isometric}\},\\ \text{Im}\mathbf{A}&=\{(\left\langle{\bm{\sigma}}\right\rangle,-\left\langle{z\bm{\sigma}}\right\rangle)\mid\bm{\sigma}\text{ is admissible}\}.\end{split}$ (35)